Exercise 1 (Adapted from Chapter 2 of Williamson text- book) 1. Imagine an economy with two firms: Firm A and Firm B. Firm A pro duces wheat and firm B produces bread. In a gven year, firm A produces 60,000 bushes of whea, sells 30,000 bushels of wheat to firm B at 85 per bushel, exports 15,000 bushes of wheat at 85 per bushel, and stores 15,000 bushels as inventory. Firm A pays 850,000 in wages to consumers. Firm B produces 50,000 loaves of bread, and sells al of it to the domestic consumers at S4 per loaf Firm B pays consumers 830,000 in wages. In addition to the 50,000 loaves of bread consumers buy from B, consumers import and consume 15,000 loaves of bread, and they pay $2 per loaf for this imported bread. Calculate Gross Domestic Product for the year us- ing (a) the product appoach, (b) the expenditure approach, and (c) the income approach
Answer #1

Under product approach gdp is calculated by aggregating the market value of all goods and services produced in an economy for an year.so computation of gdp goes like this.Apart from this goods used in manufacture of goods or rendering of services(intermediary stages) is to be excluded and only the value addition made at the next stage is counted.

Value of 60000 bushels of wheat produced by firm A= $300,000(60000*$5)

value of bread produced by firm B excluding goods used in intermediary stages

= value of finished output- value of raw materials used in production of bread

=(50000*4) - (30000*5)

=$50000

therefore gdp would be equal to $350,000(300,000 + 50000).

under expenditure approach gdp is calculated by summing up consumer spending(C),business investments in economy(I), government expenditure(G) and net exports(NX).

C= $230,000((50000 loaves *$4) + ( 15000 loaves * $2))(bread consumed by domestic consumers + consumption of imported bread)

I =$75,000(15000 bushels * $5)(wheat stored as inventory by firm A)

G= 0(there is no govt spending)

NX=

= exports -imports

=(15000 bushels * $5)(exports of wheat  by firm A) - (15000 loaves of bread * $2)(bread imports)

= $75000- $30000

= $45000

by adding all the above we get gdp as $ 350,000

under income approach gdp is measured by adding all profits received by producers of output, wages received by workers and profits received by firms or establishments.

wages paid by firm A and firm B = $ 80,000(50000+30000) (W)

Value of wheat produced by firm A - wages paid= profit earned by firm A

=> 300,000 - 50000 = $250,000

Value of bread produced by firm B - wages paid - cost of wheat purchased from firm A= profit earned by firm B

=> 200,000 - 30,000 - 150,000(30,000 Bushels * $5) = $20,000

adding the wages and profits earned by both firms gives the gdp value as

= 80,000+ 250,000 + 20,000

= $350,000.

hope i solved your question